OpenStudy (anonymous):

solve the following equation: 3-n/2=3/4

OpenStudy (anonymous):

please show me how to do this

OpenStudy (rulnick):

subtract 3 from each side: -n/2 = 3/4-3 = -2 1/4 mult by 2: -n = -4 1/2 mult by -1 n = 4 1/2 or 9/2

OpenStudy (rulnick):

... or if it is (3-n) / 2 = 3/4, then mult by 2: 3-n = 3/2 subtract 3: -n = -3/2 mult by -1: n = 3/2